OpenStudy (anonymous):

find the variation constant and an equation of variation where y varies directly as x, y=30 when x =6

jimthompson5910 (jim_thompson5910):

y=kx 30=k(6) 30/6 = k 5 = k k = 5 So the variation constant is k = 5 and the equation of variation is y = 5x

OpenStudy (anonymous):

thank you, i was lost.

jimthompson5910 (jim_thompson5910):

To show this, plug in x = 6 to get y = 5x y = 5(6) y = 30 So when x = 6, y = 30 This verifies our answer